Menlo Company distributes a single product. The company’s sales and expenses for last month follow:


Total Per Unit
Sales $ 616,000 $ 40
Variable expenses 431,200 28
Contribution margin 184,800 $ 12
Fixed expenses 144,000
Net operating income $ 40,800


Required:

1. What is the monthly break-even point in unit sales and in dollar sales?

2. Without resorting to computations, what is the total contribution margin at the break-even point?

3-a. How many units would have to be sold each month to attain a target profit of $69,600?

3-b. Verify your answer by preparing a contribution format income statement at the target sales level.

4. Refer to the original data. Compute the company's margin of safety in both dollar and percentage terms.

5. What is the company’s CM ratio? If sales increase by $83,000 per month and there is no change in fixed expenses, by how much would you expect monthly net operating income to increase?
Answer #1

Answers

  • [1]

A

Fixed expenses

$144,000

B

Contribution margin per unit

$12

C = A/B

Break even point in unit sales

                     12,000

Answer

D = C x $ 40

Break even in dollars

$480,000

Answer

  • [2]

Total contribution margin = $ 144000 which is Fixed cost because at Break even point contribution margin = Fixed cost.

  • [3]

--Part [a]

A

Target Profits

$69,600

B

Fixed expenses

$144,000

C

Contribution margin per unit

$12

D = (A+B) / C

Units to be sold

                     17,800

Answer

--Part [b]

Total

Per unit

Sales (17800 units)

$712,000

$40

Variable expenses

$498,400

$28

Contribution margin

$213,600

$12

Fixed expenses

$144,000

Net Operating Income

$69,600

  • [4]

A

Total Sales

$616,000

B

Break even in dollars

$480,000

C = A -B

Margin of Safety in dollars

$136,000

D = (C/A) x 100

MOS %

22.08%

  • [5]

A

Contribution margin per unit

$12

B

Sale price per unit

$40

C = (A/B) x 100

CM Ratio

30%

Answer

D

Increase in sales

$83,000

E = C x D

Increase in Net Operating Income

$24,900

Answer